# Language Driven Occupancy Prediction

### v1

@@ 4.2 Implementation Details
Table 4: Detailed metrics of the pseudo-labeled ground truth from different OV-Seg models on the Occ3D-nuScenes dataset. We report the mean IoU (mIoU) for semantics across different categories. The top three results among different methods are in red.
Method mIoU
@@
SAN [40] 25.53 0.00 9.40 8.47 30.16 38.00 11.43 20.20 11.39 11.54 9.55 25.73 61.86 0.40 36.41 33.00 59.42 67.10
ODISE [39] 25.80 0.00 11.45 7.83 37.66 38.10 6.94 22.74 16.28 0.52 5.58 24.59 63.09 0.27 37.16 34.12 62.72 69.47
CAT-Seg [8] 26.72 0.00 3.87 4.56 41.53 36.90 26.65 17.26 11.28 11.08 3.45 25.97 63.21 0.34 37.80 36.63 60.60 73.03
4.3 Quantitative Results
We evaluate the effectiveness of the proposed LOcc framework on three models: BEVFormer [20], BEVDet [12] and BEVDet4D [11], referred to as LOcc-BEVFormer, LOcc-BEVDet, and LOcc-BEVDet4D, respectively.
We present the quantitative results of open-vocabulary occupancy prediction in Table 1. With the input image of size $\text{256}\times\text{704}$, LOcc-BEVFormer, LOcc-BEVDet, and LOcc-BEVDet4D achieve an mIoU of 18.62, 20.29, and 22.95, respectively. Notably, LOcc-BEVDet, a more simpler model, surpasses all previous zero-shot and open-vocabulary approaches in terms of mIoU, even though many of these approaches use larger image backbones or higher-resolution inputs. Furthermore, when the image resolution is increased to $\text{512}\times\text{1408}$, the models attain an better mIoU of 19.85, 20.84, and 23.84.
4.4 Ablation Study
Label Transferring Process. Table 2 shows an analysis of label transferring process. First, we replace the segmentation maps with image features from SAN [40], requiring the learned 3D features to align with these features, thereby indirectly establishing correspondences with texts. Setting (a) demonstrates that this model results in a performance drop across all models. Next, we apply our proposed hierarchical ground truth generation pipeline using single-frame vocabularies. As shown in setting (b), the model achieves mIoU scores of 17.87, 19.61, and 22.15, respectively. Finally, by treating multiple consecutive frames as a unified sequence and consolidating the vocabulary from each frame into a cohesive set for pseudo-label assignment, as indicated in setting (c), we observe additional performance gains, further validating the effectiveness of our voxel-to-language modeling approach.
Table 5: Ablation study on the OV-Seg models.
@@ 4.4 Ablation Study
Scene Reconstruction Process. Voxel-based model-view projection treats all the voxels as surface ones, overlooking the occlusion problem and leading to noisy and sparse voxel-to-text correspondences. Instead, each frame of LiDAR is easily calibrated with corresponding images during the data collection process, with much less misprojections. Setting (e) demonstrates that using ground truth from voxel-based model-view projection to train the OVO models results in performance drops to 13.77, 14.49, and 15.41. In contrast, performing scene reconstruction with nearest-point voxelization yields notable improvements, as shown in setting (d). In contrast, majority-voting voxelization can bring better performance, where the mIoU is further improved to 18.62, 20.29, and 22.68 mIoU, as illustrated in setting (c). The results prove that majority-voting is more robust and less susceptible to the influence of single-point projection errors.
To further demonstrate the effectiveness of our transitive semantic labeling pipeline, we present detailed metrics for the pseudo-labeled ground truth generated under various settings in Table 3. Consistent with prior analyses, our pipeline with majority-voting voxelization, yields the highest-quality ground truth in terms of mIoU. Additionally, these results confirm that improvements in the generated pseudo-labeled ground truth directly enhance the performance of the final OVO model.
@@
OV-Seg Models. Table 4 lists the metric values of the pseudo-labeled ground truth over the human-labeled ground truth using different OV-Seg models: SAN [40], ODISE [39] and CAT-Seg [8]. As shown, all models perform well on larger objects but focus on different smaller objects. Then we use the generated pseudo-labeled ground truth to train the OVO networks, with the results listed in Table 5. The networks achieve comparable performance with different segmentation models. In this work, we use SAN [40] for fair comparison with previous approaches.
Language Autoencoder. We conduct experiments to verify the effectiveness of the autoencoder. Without it, the model is required to generate 3D features with a dimension of 512. In this case, we set the voxel feature dimension after the 2D-to-3D transformation for BEVDet [12] and BEVDet4D [11] to 128. For BEVFormer [20], the BEV queries are kept at a dimension of 256, as using a larger dimension exceeds the GPU memory. As shown in Table 6, introducing autoencoder can effectively reduce the memory requirements with better performance. Here, we hypothesize that this is because CLIP model is trained using 400 million (image, text) pairs, thus its high-dimensional space could be highly compact. Instead, the number of extracted texts is around one thousand in this work, which is significantly smaller than the number of texts used in CLIP training, resulting a sparse space and allowing us to further compress it.
@@ 4.5 Qualitative Results
Table 6: Ablation study on the language autoencoder.
Method w/o autoencoder w/ autoencoder
[…]
LOcc-BEVFormer 17.36 1.41 61.80 18.62 1.18 61.70
LOcc-BEVDet 20.25 2.22 92.40 20.29 1.11 41.60
LOcc-BEVDet4D 22.62 2.33 100.00 22.68 1.11 48.00

### v3

@@ 4.2 Evaluation Benchmarks
a
4.3 Implementation Details
[…]
Table 4: Open-vocabulary occupancy prediction results on the Occ3D-nuScenes dataset. We report the mean IoU (mIoU) for semantics across different categories, along with per-class semantic IoUs. The best results among the methods are in bold.
Model Image Backbone Image Size mIoU
$\blacksquare$ bicycle
$\blacksquare$ motorcycle
$\blacksquare$ traffic cone
$\blacksquare$ sidewalk
@@ 4.3 Implementation Details
LOcc-BEVFormer (ours) ResNet-101 256 $\times$ 704 23.26 18.71 22.21 24.33 48.14 0.00 14.48 28.16 33.88 4.32 9.19 2.37 22.45 72.09 0.00 34.32 31.22 29.58
LOcc-BEVDet (ours) ResNet-50 256 $\times$ 704 23.50 12.21 17.33 17.24 47.03 0.00 13.43 32.67 35.54 8.68 10.01 2.54 27.33 72.16 0.00 36.31 34.47 32.57
LOcc-BEVDet4D (ours) ResNet-50 256 $\times$ 704 26.91 20.34 23.77 23.93 51.47 0.00 15.63 29.54 39.89 10.99 14.01 6.21 31.72 73.23 0.00 40.11 39.00 37.76
Table 5: Ablation study on the OV-Seg models. The best results among different methods are in bold.
@@
4.4 Zero-shot Occupancy Prediction
We evaluate the effectiveness of the proposed LOcc framework on three models: BEVFormer [23], BEVDet [12] and BEVDet4D [11], referred to as LOcc-BEVFormer, LOcc-BEVDet, and LOcc-BEVDet4D, respectively. The comparison methods includes those that rely solely on images [14, 54] and those that utilize unlabeled LiDAR data to assist in geometry prediction [35, 56]. To ensure a fair comparison with POP-3D, which uses single-frame LiDAR for geometry supervision, we replace the single-frame LiDAR data with dense binary occupancy ground truth [56]. For OV-Seg models, we replace LSeg [18] with more powerful SAN [44].
[…]
4.5 Ablation Study
Label Transferring Process. Table 2 provides an analysis of the label transferring process. First, we replace the segmentation maps with image features extracted from SAN [44], requiring the learned 3D features to align with these image features. This configuration indirectly establishes correspondences with textual labels. Setting (a) demonstrates that it leads to a performance drop in the generated pseudo-labeled ground truth, which in turn degrades the performance of all models. Next, we apply our proposed semantic transitive labeling pipeline using only single-frame vocabularies. As shown in setting (b), the mIoU of the pseudo-labeled ground truth increases to 22.66, and the metrics for all models improve to 17.87, 19.61, and 22.15, respectively. Finally, by treating multiple consecutive frames as a unified sequence and consolidating the vocabulary from each frame into a cohesive set for pseudo-label assignment, as indicated in setting (c), we observe additional performance gains, further validating the effectiveness of our label transferring process.
@@ 4.5 Ablation Study
Scene Reconstruction Process. The voxel-based model-view projection treats all voxels as point clouds, making it challenging to determine whether a voxel lies on the surface after projection. This coarse approach overlooks the occlusion problem, leading to noisy voxel-to-text correspondences. Setting (e) demonstrates that the pseudo-labeled ground truth from voxel-based model-view projection can only achieve an mIoU of 19.55. Using this unsatisfied ground truth to train the OVO models results in a huge performance drop to 13.77, 14.49, and 15.41. Instead, each frame of LiDAR can be accurately calibrated with corresponding images during the data collection process, significantly reducing misprojections. Thus, performing scene reconstruction with nearest-point voxelization yields notable improvements, as shown in setting (d). Furthermore, majority-voting voxelization can mitigate the issue of wrong segmentation results, achieving better results with the mIoU of pseudo-labeled and models improving to 25.53, 18.62, 20.29, and 22.68, as illustrated in setting (c). These results prove that majority-voting is more robust and less susceptible to the influence of single-point projection errors. The overall ablation results prove that the quality of the pseudo-labeled ground truth highly influences the performance of OVO models. In Table 3, we present detailed metrics for the pseudo-labeled ground truth generated under various settings. Consistent with prior analyses, our pipeline yields better results on most of categories, further confirming the effectiveness of our proposed pipeline.
@@
OV-Seg Models. Table 5 lists the results of the OVO networks trained on the pseudo-labeled ground truth generated from different OV-Seg models: SAN [44], ODISE [43] and CAT-Seg [8]. In this work, we use SAN [44] for fair comparison with previous approaches [56].
4.6 Open-vocabulary Occupancy Prediction
Table 4 lists the detailed metrics of open vocabulary occupancy prediction. We use bicycle, motorcycle, traffic cone, and sidewalk as base classes, while treating the remaining categories as novel classes. With more seen classes, the overall performance for the models rise. Besides, the IoUs on unseen classes are always competitive, proving the open-vocabulary capability of our methods.
@@ 5 Conclusions
Acknowledgments
We thank the reviewers for the valuable discussions. This research was supported by the National Key Research and Development Program of China under grant 2023YFB3209800, in part by the National Natural Science Foundation of China under grant 62301484, in part by the Ningbo Natural Science Foundation of China under grant 2024J454, in part by the Aeronautical Science Foundation of China under grant 2024M071076001 and in part by the Zhejiang Provincial Natural Science Foundation of China under Grant No. LD24F030001.